OpenStudy (anonymous):

the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (cgreenwade2000):

When it asks for the equation of a line, you should instantly think of SLOPE INTERCEPT FORM. y=mx+b Here, you have the slope and an x and y value.

OpenStudy (anonymous):

substitute m for -1/3 then substitute 6 for x and -4 for y and solve for b. when your done your equation should look something like this y=5x+6

OpenStudy (cgreenwade2000):

y=5x+6 isn't the answer, to clarify.

OpenStudy (anonymous):

-4= -1/3(6) +b -4=-2+b b=-2 Answer: y=-1/3x-2

OpenStudy (anonymous):

yup lol